# Characterization of NBI-driven shear Alfvén waves in the TJ-II stellarator using Mirnov probes and electrostatic potential fluctuation measurements

### v1

@@ 1 Introduction
The main objective of the experiments described in this paper was to excite AEs in different scenarios and study their spatial structure, focusing on the mode number identification () by means of poloidal and helical arrangements of Mirnov coils [ascasibarMeasurementsMagnetic2022] on the one hand and, on the other hand, on the measurement of the radial profile of perturbed potential using Heavy Ion Beam Probes (HIBP). The measurement of the poloidal mode structure is routinely done since the poloidal array of Mirnov coils was commissioned and the determination of the poloidal number has been employed in previous studies [jimenez-gomezAlfvenEigenmodes2011, vanmilligenMHDMode2011]. Only very recently we have added the possibility to determine the toroidal mode number [ascasibarMeasurementsMagnetic2022] and the polarization of the wave detected at the coils locations. Beyond the interest related to model validation, the difficulty of this measurement, in particular in non-axisymmetric configurations, calls for a complete experimental study where we know that we are varying the spectrum of possible modes and can thus confirm that these variations are reflected in the helical array data and the toroidal mode numbers reconstructed from this data. Both measurements, magnetics and HIBP, along with the recent development of a synthetic diagnostic for Mirnov coil arrays [pons-villalongaExploringOperational2024] allow us to establish a criterion for the validity of the mode analysis results.
@@
On-axis Electron Cyclotron Current Drive (ECCD) was chosen as an actuator to modify the spectrum of AEs due to its minimal impact on the plasma density and temperature profiles compared to the effect it has on the plasma current, which isolates the rotational transform profile as the driver of changes in the shear Alfvén wave (SAW) spectrum [cappaStabilityAnalysis2021]. To obtain a comprehensive database of different excited AEs, we conducted a parameter scan on magnetic configuration and EC beam injection angle. An essential piece of information when performing validation tasks is the rotational transform profile. Unfortunately, as we will show here, the changes induced by ECCD or Neutral Beam Current Drive (NBCD), which are important from the point of view of their impact on the Alfvén continuum, are not intense enough to be measurable by iota diagnostics such as Motional Stark Effect (MSE) [mccarthySpectrallyResolved2015]. Therefore we must rely on an estimation of the different external current sources and simulate the time evolution of rotational transform by calculating the radial diffusion of the shielding current.
@@
This paper is organized as follows: section 2 contains a description of the experimental set-up and the relevant plasma parameters that were achieved in the experiments. Then, in section 3 we describe the mode analysis technique and the whole set of measured mode numbers. Next, in section 4, after a brief description of the HIBP measurements, we present several examples of the perturbed electric potential profiles associated with the modes. In section 5, we solve the shielding current radial diffusion equation in cylindrical geometry taking the total experimental current as boundary condition and approximate profiles for the external current sources. This allows us to estimate the evolution of the rotational transform profile. A short discussion on the difficulty of resolving changes in using MSE measurements is also included in this section. Although the main purpose of the paper is the experimental characterization of instabilities in different scenarios, in section 6 we will take a couple of cases and compare the observations with the Alvén continuum calculated with the STELLGAP code [spongShearAlfven2003]. Finally, we come to our conclusions and lay the future groundwork for the implementation of the validation studies.
@@ 2 Experimental arrangement and results
To investigate the spatial structure of the AEs destabilized in different magnetic configurations and with variable on-axis ECCD settings, the following experimental set-up was used. The heating sequence is shown in figure 2. The plasma is started using ECRH, with both beams switched-on simultaneously. One of them is set to inject perpendicularly to the magnetic axis (ECRH2) and the other (ECRH1) is launched with different incidence angles that are varied on a shot to shot basis. After ms, NBI2 is switched on, creating a fast-particle population that heats the plasma when slowing down and at the same time induces current (NBCD) and drives the AEs. The ECRH2 beam with perpendicular incidence is turned off shortly after the injection of NBI, while the ECRH1 beam remains on to both drive ECCD and help to maintain a constant plasma density during the rest of the shot.
We have performed the same scan using three distinct magnetic configurations with different values of rotational transform () on-axis and very similar low shear. For our present purposes, although the central rotational transform can be varied between and , we will refer to them as low (), medium (or standard, ) and high iota () configurations. The vacuum profiles are represented in figure 3, along with a poloidal cross section of the magnetic surfaces of each configuration at .
@@
Negative values of correspond to co-ECCD, generating positive current (increasing ), while positive ones stand for counter-ECCD, that generates negative current (decreasing ). Table 1 shows the ECRH injection angle in degrees for each value of .
@@
A good density control is exhibited during the time interval of interest (between and ms) for the low and medium configurations. This simplifies the mode number analysis, allows for reproducible measurements of mode potential radial structure and is fundamental when carrying out validation studies, that benefit from having a steady state fast ion distribution and a steady state NBI-driven current source. In the case of the high configuration, density control was harder to achieve due to degradation of wall condition, requiring an increase in ECRH power during NBI phase (up to 200 kW) and also a decrease in pulse length.
@@ 3 Magnetic fluctuations and mode number measurements
There are two arrays of Mirnov coils, that digitize the magnetic fluctuations at a rate of 1 MS/s. One of them, hereinafter referred to as poloidal array, consists of 25 single-axis coils distributed in a vertical plane at constant toroidal angle, and provide information on the poloidal mode number by measuring magnetic fluctuations in the (approximate) poloidal direction [jimenez-gomezAlfvenEigenmodes2011, jimenez-gomezAnalysisMagnetohydrodynamic2007]. The other array, hereinafter referred to as helical array, is distributed toroidally along a full period of the device, and is comprised of two up-down symmetrical sub-arrays of 32 tri-axial sensors each [ascasibarMeasurementsMagnetic2022] that follow the helical path of the plasma. These arrays are used to identify both toroidal () and poloidal () numbers of the excited modes. Additionally, there are also two vertical arrays located at different toroidal angles not used for mode periodicity measurements. The mode analysis is carried out using the Lomb periodogram [lombLeastsquaresFrequency1976, zegenhagenAnalysisAlfven2006], that can handle non-equispaced data points. An in-depth discussion on the mode analysis can be found in [pons-villalongaExploringOperational2024], so here it will just be outlined. For the mode analysis, each coil has to be mapped to a pair of magnetic angles, that in our case will be the Boozer angles of the closest point to the coil in the Last Closed Flux Surface (LCFS). These angles are found using the vacuum equilibrium of each configuration, as tests with different plasma current profiles show that their variation is negligible in comparison with the error induced by the uncertainty in the positioning of the coils. Figure 7 shows the distribution of the poloidal and helical arrays around the plasma and the Boozer angles for the standard configuration.
@@ 3.1 Observed eigenmodes
The spectrum of NBI driven modes observed in each of the three magnetic configurations, for all the values of explored, is shown in figure 8. The fact that the plasma density is not exactly constant throughout the shot allows us to identify the Alfvénic nature of the excited modes by their slow evolution in frequency, an effect which is largely amplified when the ECRH heating shutdown at ms produces a sudden increase in plasma density and a concomitant reduction in mode frequency ( as expected for AEs). This figure also illustrates how the spectrum of modes is modified as the ECRH1 beam is steered along the magnetic axis. In the medium case, the shear Alfvén spectrum clearly evolves from a situation with three main branches well separated in frequency to a situation in which only a single intense mode is detected. Similar changes are observed in the high case while almost no modes are observed for the low configuration during the ECRH+NBI phase. Low frequency MHD modes are present in all shots, appearing less clearly in the low configuration. This set of shots shows a wide array of behaviours of the magnetic fluctuations. Among the most remarkable we find:
Modes that emerge and vanish, for example in shots 53491 and 53487, where transient modes can be observed above the kHz mode that is present for the whole shot. This could be caused by the strong low frequency inductive oscillations in current mentioned above and the associated changes in iota or by non-linear energy transfer between the modes, both effects being outside the scope of the paper.
@@
An example of chirping mode is shown in figure 9. The mode displays rapid chirping behaviour, with a repetition frequency of about 1 kHz. These chirping modes are more difficult to analyze in terms of mode numbers, as the rapid sweep in frequency deteriorates the precision of the analysis technique, that is discussed below. Aditionally, figure 10 shows a comparison of spectrograms obtained via Short Time Fourier Transforms (STFT) and the DMUSIC method [kleiberModernMethods2021] for shot 53490. A closer look tells us that what looks like a single intense mode is actually formed by three distinct frequency components separated by about 4 kHz from each other, and that remain remarkably stable during the considered time interval. The DMUSIC method is very useful in these cases, as it helps in the precise determination of the mode frequency, that is needed for the periodogram. Figure 10 also shows the bandpass filtered signal (between 250 and 310 kHz), that shows large amplitude oscillations interleaved with periods
of lower amplitude as corresponds to a combination of three near-frequency modes of similar amplitude. Despite the clear frequency structure, the mode analysis cannot resolve different mode numbers for each of the observed frequencies. Even applying very narrow bandpass filter intended to isolate each frequency contribution the mode number result remains the same (see 3.2).
@@ 3.2 Mode number identification
The Lomb periodogram of a signal , where is a time index and a coil index, is
where is the phase term, is a normalization factor, and is a phase shift term given by
@@
The Lomb periodogram requires as input the frequency of the mode, that is obtained from the spectrogram, and the integer pair of mode numbers (), to which from now on we will refer to as , over which a scan is carried out to obtain a two-dimensional distribution of intensities. Given the direction chosen for the coil distribution arrangement and the TJ-II boozer coordinate system, the sign criterion (+) in the phase term is such that a mode number with positive and positive propagates toroidally along the direction of the magnetic field and poloidally along the direction of the ion diamagnetic drift.
@@
After determining the real orientation, and correct phase sign of each set of coils by means of dedicated calibration experiments described in [ascasibarMeasurementsMagnetic2022], the measured signals can be projected over any desired polarization vector basis. In this work we will use the projection along the normalized tangent basis vector , also in Boozer coordinates, as it proved to be the best-performing basis for the mode number analysis [pons-villalongaExploringOperational2024]. A suitable collection of time and frequency intervals are selected for each mode, and the signals, calibrated and expressed in the chosen polarization basis, are then filtered and normalized to minimize distance effects on the periodogram result.
@@
The signals from the poloidal and helical arrays cannot be included in the same periodogram, as a difference in coil orientation induces a phase shift in the signal that compromises the accuracy of the analysis. To work around that, periodograms for the poloidal and helical arrays are computed independently and then multiplied. Furthermore, this method also helps work around the main limitation of the helical array, that is, that the coils follow a straight path in Boozer coordinates so, in an aliasing-like phenomenon, a strip of modes with slope appears that contains the true pair of mode numbers. The product with the periodogram of the poloidal array, that suffers to a much lesser extent from the same issue, enables in principle the determination of the true pair.
@@
Figure 11 shows an example of this process applied to the a low frequency mode observed in one of the medium configuration shots, with the periodograms for the poloidal array (left), helical array (center) and their product (right). The identified pair of mode numbers is . As we have already mentioned, a fundamental drawback due to the non-axisymmetric configuration and the spatial distribution of the coils is that there are still modes with different pairs of characteristic mode numbers that appear with almost the same probability in the analysis and we must use arguments based on the expected values of the rotational transform to discard some combinations of modes versus others. Low frequency MHD modes have shown to be easier to identify than AEs with higher frequency even though they exhibit larger bandwith, as is the case of the modes analyzed in this work. The quality of the mode analysis deteriorates with rapid changes in frequency and mode intensity. Furthermore, systematic errors during data acquisition and reduced coil availability cause a broadening of the mode strips and sometimes introduce false positives. High frequency broadband modes are also difficult to analyze using this method, and require more sophisticated techniques such as the Stochastic System Identification (SSI) [kleiberModernMethods2021]. Especially deleterious for the analysis are spurious phase shifts [horvathReducingSystematic2015], that are also difficult to characterize experimentally for the coil arrays of TJ-II.
@@ 4 HIBP Measurements
Figure 17 shows an example of the measured quantities for the case in the medium configuration (shot #53492). The spectrogram of the magnetic fluctuations corresponding to a shot with the same conditions (#53491) is shown in figure 8. The more or less stationary plasma, which translates into a constant mode frequency, allows us to average over repetitive scans of the heavy ion beam and retrieve a radial profile of the mode. This result, that can already be guessed by inspecting the spectrogram of plasma potential fluctuations presented in 17 is shown in figure 18 again for .
In this case, the spectrogram of magnetic fluctuations shows three distinct modes around 180, 215 and 310 kHz. The one with lower frequency is barely seen by HIBP1 in either density or potential fluctuations. The radial profile measurements reveal a more core localized 215 kHz mode with an approximately symmetrical profile while the 310 kHz mode is shifted towards the medium region of the plasma and exhibits of somewhat asymmetrical profile. The corresponding mode numbers extracted from the Mirnov coils data were given in figure 15 and appear also in the figures. The bottom panel in figure 18 shows the dependence of the secondary beam intensity (Cs++) on the radial position. As shown in reference [khabanovDensityProfile2019], the density profile, roughly proportional to the beam intensity profile can be reconstructed from this measurement. Note that, consistently with the profiles shown in figure 6, the intensity gradient region falls between and . This is important when considering the excitation of low frequency MHD modes and its relation to the rotational transform profile that will be explored in section 5.
As mentioned above, for some modes a radial asymmetry in the potential profile is clearly observed. This is consistent with the fact that these are probably gap modes [kolesnichenkoAlfvenEigenmodes2002]. These modes consist of coupled modes with different ’s and ’s that evolve together. The difference in poloidal mode numbers produces the asymmetry in the potential profiles. On the other hand, as it was demonstrated in reference [pons-villalongaExploringOperational2024] using a synthetic Mirnov coils diagnostic, in the case of coupled modes, the magnetic coils cannot resolve different ’s when the difference between them is or . In these case, only the dominant mode numbers result from the analysis. The radial profiles of the mode potential have been measured for several values of and additional results for and are shown in figures 19, and 20, 21 respectively.

### v2

@@ 1 Introduction
The main objective of the experiments described in this paper was to excite AEs in different scenarios and study their spatial structure, focusing on the mode number identification () by means of poloidal and helical arrangements of Mirnov coils [ascasibarMeasurementsMagnetic2022] on the one hand and, on the other hand, on the measurement of the radial profile of perturbed potential using Heavy Ion Beam Probes (HIBP). The measurement of the poloidal mode structure is routinely done since the poloidal array of Mirnov coils was commissioned and the determination of the poloidal number has been employed in previous studies [jimenez-gomezAlfvenEigenmodes2011, vanmilligenMHDMode2011]. Only very recently we have added the possibility to determine the toroidal mode number [ascasibarMeasurementsMagnetic2022] and the polarization of the wave detected at the coils locations. Beyond the interest related to model validation, the difficulty of this measurement, in particular in non-axisymmetric configurations, calls for a complete experimental study where we know that we are varying the spectrum of possible modes and can thus confirm that these variations are reflected in the helical array data and the toroidal mode numbers reconstructed from this data. Both measurements, magnetics and HIBP, along with the recent development of a synthetic diagnostic for Mirnov coil arrays [pons-villalongaExploringOperational2024] allow us to establish appropriate criteria to support the validity of the mode analysis results.
@@
On-axis Electron Cyclotron Current Drive (ECCD) was chosen as actuator to modify the spectrum of AEs due to its minimal impact on the plasma density and temperature profiles compared to the effect it has on the plasma current, which isolates the rotational transform profile () as the driver of changes in the shear Alfvén wave (SAW) spectrum [cappaStabilityAnalysis2021]. To obtain a comprehensive database of different excited AEs, we conducted a parameter scan on magnetic configuration and EC beam injection angle. An essential piece of information when performing validation tasks is the rotational transform profile. Unfortunately, as we will show here, the changes induced by ECCD or Neutral Beam Current Drive (NBCD), which are important from the point of view of their impact on the Alfvén continuum, are not intense enough to be measurable by iota diagnostics such as Motional Stark Effect (MSE) [mccarthySpectrallyResolved2015]. Therefore we must rely on an estimation of the different external current sources and simulate the time evolution of rotational transform by calculating the radial diffusion of the shielding current.
@@
This paper is organized as follows: section 2 contains a description of the experimental set-up and the relevant plasma parameters that were achieved in the experiments. Then, in section 3 we describe the mode analysis technique and the whole set of measured mode numbers. Next, in section 4, after a brief description of the HIBP measurements, we present several examples of the perturbed electrostatic potential profiles associated with the modes. In section 5, we solve the shielding current radial diffusion equation in cylindrical geometry taking the total experimental current as boundary condition and approximate profiles for the external current sources. This allows us to estimate the evolution of the rotational transform profile. A short discussion on the difficulty of resolving changes in using MSE measurements is also included in this section. Although the main purpose of the paper is the experimental characterization of instabilities in different scenarios, in section 6 we will take a couple of cases and compare the observations with the Alvén continuum calculated with the STELLGAP code [spongShearAlfven2003]. Finally, we come to our conclusions and lay the future groundwork for the implementation of the validation studies.
@@ 2 Experimental arrangement and results
To investigate the spatial structure of the AEs destabilized in different magnetic configurations and with variable on-axis ECCD settings, the following experimental set-up was used. The heating sequence is shown in figure 2. The plasma is started using ECRH, with both beams switched-on simultaneously. One of them is set to inject perpendicularly to the magnetic axis (ECRH2) and the other (ECRH1) is launched with different incidence angles that are varied on a shot to shot basis. After ms, NBI2 is switched on, creating a fast-particle population that heats the plasma when slowing down and at the same time induces current (NBCD) and drives the AEs. The ECRH2 beam with perpendicular incidence is turned off shortly after the injection of NBI2, while the ECRH1 beam remains on to both drive ECCD and help to maintain a constant plasma density during the rest of the shot.
We have performed the same scan using three distinct magnetic configurations with different values of rotational transform on-axis and very similar low shear. For our present purposes, although the central rotational transform can be varied between and , we will refer to them as low (), medium (or standard, ) and high iota () configurations. The vacuum profiles are represented in figure 3, along with a poloidal cross section of the magnetic surfaces of each configuration at .
@@
Negative values of correspond to co-ECCD, generating positive current (increasing ), while positive ones stand for counter-ECCD, that generates negative current (decreasing ). Table 1 shows the ECRH1 injection angle in degrees for each value of .
@@
A good density control is exhibited during the time interval of interest (between and ms) for the low and medium configurations. This simplifies the mode number analysis, allows for reproducible measurements of mode potential radial structure and is fundamental when carrying out validation studies, that benefit from having a steady state fast ion distribution and a steady state NBI2-driven current source. In the case of the high configuration, density control was harder to achieve due to degradation of wall condition, requiring an increase in ECRH1 power during NBI2 phase (up to 200 kW) and also a decrease in pulse length.
@@ 3 Magnetic fluctuations and mode number measurements
There are two arrays of Mirnov coils, that digitize the magnetic fluctuations at a rate of 1 MS/s. One of them, hereinafter referred to as poloidal array, consists of 25 single-axis coils distributed in a vertical plane at constant toroidal angle, and provide information on the poloidal mode number by measuring magnetic fluctuations in the (approximate) poloidal direction [jimenez-gomezAlfvenEigenmodes2011, jimenez-gomezAnalysisMagnetohydrodynamic2007]. The other array, hereinafter referred to as helical array, is distributed toroidally along a full period of the device, and is comprised of two up-down symmetrical sub-arrays of 32 tri-axial sensors each [ascasibarMeasurementsMagnetic2022] that follow the helical path of the plasma. These arrays are used to identify both toroidal () and poloidal () numbers of the excited modes. Additionally, there are also two vertical arrays located at different toroidal angles not used for mode periodicity measurements. The mode analysis is carried out using the Lomb periodogram [lombLeastsquaresFrequency1976, zegenhagenAnalysisAlfven2006], that can handle non-equispaced data points. An in-depth discussion on the mode analysis can be found in [pons-villalongaExploringOperational2024], so here it will just be outlined. For the mode analysis, each coil has to be mapped to a pair of magnetic angles, that in our case will be the Boozer angles of the closest point to the coil in the Last Closed Flux Surface (LCFS). These angles are found using the vacuum equilibrium of each configuration, as tests with VMEC equilibria calculated with different plasma current profiles show that their variation is negligible in comparison with the error induced by the uncertainty in the positioning of the coils. Figure 7 shows the distribution of the poloidal and helical arrays around the plasma and the Boozer angles for the standard configuration.
@@ 3.1 Observed eigenmodes
The spectrum of NBI2 driven modes observed in each of the three magnetic configurations, for all the values of explored, is shown in figure 8. The fact that the plasma density is not exactly constant throughout the shot allows us to identify the Alfvénic nature of the excited modes by their slow evolution in frequency, an effect which is largely amplified when the ECRH1 heating shutdown at ms produces a sudden increase in plasma density and a concomitant reduction in mode frequency ( as expected for AEs). This figure also illustrates how the spectrum of modes is modified as the ECRH1 beam is steered along the magnetic axis. In the medium case, the shear Alfvén spectrum clearly evolves from a situation with three main branches well separated in frequency to a situation in which only a single intense mode is detected. Similar changes are observed in the high case while almost no modes are observed for the low configuration during the ECRH1+NBI2 phase. This set of shots shows a wide array of behaviours of the magnetic fluctuations. Among the most remarkable we find:
Modes that emerge and vanish, for example in shots 53491 and 53487, where transient modes can be observed above the kHz mode that is present during the entire NBI2 phase. This could be caused by the strong low frequency inductive oscillations in current mentioned above and the associated changes in iota or by non-linear energy transfer between the modes, both effects being outside the scope of the paper.
@@
An example of chirping mode, analyzed via Short-Time Fourier Transforms (STFT) and the DMUSIC method [kleiberModernMethods2021] is shown in figure 9. The mode displays rapid chirping behaviour, with a repetition frequency of about 1 kHz. These chirping modes are more difficult to analyze in terms of mode numbers, as the rapid sweep in frequency deteriorates the precision of the analysis technique discussed below. Aditionally, figure 10 shows again a comparison of STFT and DMUSIC spectrograms calculated for shot 53490. A closer look tells us that what looks like a single intense mode is actually formed by three distinct frequency components separated by about 4 kHz from each other, and that remain remarkably stable during the considered time interval. The DMUSIC method is very useful in these cases, as it helps in the precise determination of the mode frequency, that is needed for the periodogram. Figure 10 also shows the bandpass filtered signal (between 250 and 310 kHz), that shows large amplitude oscillations interleaved with periods
of lower amplitude as corresponds to a combination of three near-frequency modes of similar amplitude. Despite the clear frequency structure, the analysis cannot resolve different mode numbers for each of the observed frequencies. Even applying very narrow bandpass filter intended to isolate each frequency contribution the mode number result remains the same (see section 3.2).
@@ 3.2 Mode number identification
The Lomb periodogram of a set of signals , where is a time index and a coil index, is given by
where is the phase term, is a normalization factor, and is a phase shift term, given by
@@
The Lomb periodogram requires as input the frequency of the mode, that is obtained from the spectrogram, and the integer pair of mode numbers (), noted to as hereinafter, over which a scan is carried out to obtain a two-dimensional distribution of intensities. Given the direction chosen for the coil distribution arrangement and the TJ-II boozer coordinate system, the sign criterion (+) in the phase term is such that a mode number with positive and positive propagates toroidally along the direction of the magnetic field and poloidally along the direction of the ion diamagnetic drift.
@@
After determining the real orientation, and correct phase sign of each set of coils by means of dedicated calibration experiments described in [ascasibarMeasurementsMagnetic2022], the measured can be projected over any desired polarization vector basis (only in the case of the helical array). In this work we will use the projection of along the normalized tangent basis vector , also in Boozer coordinates, as it proved to be the best-performing basis for the mode number analysis [pons-villalongaExploringOperational2024]. A suitable collection of time and frequency intervals are selected for each mode, and the signals, calibrated and expressed in the chosen polarization basis, are then filtered and normalized to minimize distance effects on the periodogram result.
@@
The signals from the poloidal and helical arrays cannot be included in the same periodogram, as a difference in coil orientation induces a phase shift in the signal that compromises the accuracy of the analysis. To work around that, periodograms for the poloidal and helical arrays are computed independently and then multiplied. Furthermore, this method also helps work around the main limitation of the helical array, that is, that the coils follow a straight path in Boozer coordinates so, in an aliasing-like phenomenon, a strip of pairs of mode numbers with slope , that contains the true pair of mode numbers, appears in the periodogram. The product with the periodogram of the poloidal array, that suffers to a much lesser extent from the same issue, enables in principle the determination of the true pair.
@@
Figure 11 shows an example of this process applied to a low frequency mode observed in one of the medium configuration shots, with the periodograms for the poloidal array (left), helical array (center) and their product (right). The identified pair of mode numbers is . As we have already mentioned, a fundamental drawback due to the non-axisymmetric configuration and the spatial distribution of the coils is that there are still modes with different pairs of characteristic mode numbers that appear with almost the same probability in the analysis and we must use arguments based on the expected values of the rotational transform to discard some combinations of modes versus others. Low frequency MHD modes have shown to be easier to identify than AEs with higher frequency even though they exhibit larger bandwith, as is the case of the modes analyzed in this work. The quality of the mode analysis deteriorates with rapid changes in frequency and mode intensity. Furthermore, systematic errors during data acquisition and reduced coil availability cause a broadening of the mode number strips and sometimes introduce false positives. High frequency broadband modes are also difficult to analyze using this method, and require more sophisticated techniques such as the Stochastic System Identification (SSI) [kleiberModernMethods2021]. Especially deleterious for the analysis are spurious phase shifts [horvathReducingSystematic2015], that are also difficult to characterize experimentally for the coil arrays of TJ-II.
@@ 4 HIBP Measurements
Figure 17 shows an example of the measured quantities for the case in the medium configuration (shot #53492). The spectrogram of the magnetic fluctuations corresponding to a shot with the same conditions (#53491) is shown in figure 8. The more or less stationary plasma, which translates into a constant mode frequency, allows us to average over repetitive scans of the heavy ion beam and retrieve a radial profile of the mode. This result, that can already be guessed by inspecting the spectrogram of plasma potential fluctuations presented in figure 17 is shown in figure 18.
In this case, the spectrogram of magnetic fluctuations shows three distinct modes around 180, 215 and 310 kHz. The one with lower frequency is barely seen by HIBP1 in either density or potential fluctuations. The radial profile measurements reveal a more core localized 215 kHz mode with an approximately symmetrical profile while the 310 kHz mode is shifted towards the medium region of the plasma and exhibits of somewhat asymmetrical profile. The corresponding mode numbers extracted from the Mirnov coils data were given in figure 15 and appear also in the figures. The bottom panel in figure 18 shows the dependence of the secondary beam intensity (Cs++) on the radial position. As shown in reference [khabanovDensityProfile2019], the density profile, roughly proportional to the beam intensity profile can be reconstructed from this measurement. Note that, consistently with the density profiles shown in figure 6, the intensity gradient region falls between and . This is important when considering the excitation of low frequency MHD modes and its relation to the rotational transform profile that will be explored in section 5.
As mentioned above, for some modes a radial asymmetry in the potential profile is clearly observed. This is consistent with the fact that these are probably gap modes [kolesnichenkoAlfvenEigenmodes2002]. These modes consist of coupled modes with different ’s and ’s that evolve together. The difference in poloidal mode numbers produces the asymmetry in the potential profiles. On the other hand, as it was demonstrated in reference [pons-villalongaExploringOperational2024] using a synthetic Mirnov coils diagnostic, in the case of coupled modes, the magnetic coils cannot resolve different ’s when the difference between them is or . In these case, only the dominant mode numbers result from the analysis. The radial profiles of the mode potential have been measured for several values of and additional results for and for the other magnetic configurations are shown in figures 19, and 20, 21 respectively.